# Redefining Machine Unlearning: A Conformal Prediction-Motivated Approach

### v1

@@ 2.3.2 Confidence Level and Calibration Data
In conformal prediction, miscoverage rate $\alpha$ and calibration set size are two crucial parameters for determining prediction set coverage and reliability. We next discuss the suitable settings for these two parameters and the rationale behind our chosen configuration.
In terms of $\alpha$, a higher confidence level (which is $1-\alpha$) provides more reliable coverage but typically results in larger prediction sets. Common settings of confidence level are $0.95$ and $0.9$, corresponding to $\alpha$ values of $0.05$ and $0.1$, respectively. To make our work more broadly applicable and provide a comprehensive reference, we expand our experiments by setting $\alpha$ to a wider range of values: $0.05$, $0.1$, $0.15$, and $0.2$. Notably, our default and recommended value for $\bm{\alpha}$ is $\bm{0.1}$ and subsequent experimental analysis is based on $\bm{\alpha=0.1}$. Higher values of $\alpha$ would result in smaller, potentially more precise prediction sets, but this reduction comes at the cost of decreased confidence, which may not be acceptable in many applications.
Table 4: Unlearning performance of 9 unlearning methods on CIFAR-10 with ResNet-18 in $10\%$ random data forgetting scenario. The sign $\uparrow$ represents the greater is better, while $\downarrow$ denotes ideally small. The results are average values from $3$ independent trials and the standard deviation values are reported in Appendix B. The performance gap relative to the RT method is represented in () in results tables. It shows the unlearning methods that excel under traditional metrics (UA, RA and TA) do not necessarily perform well under the CR metric.
Coverage Set Size CR
RT 0.1 $0.881({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.895({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.934({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.947({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.943({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.945({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.192$
[…]
FT 0.1 $0.968({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.087})$ $0.899({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $0.969({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.035})$ $0.924({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.023})$ $0.998({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.972({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$ $0.079$
[…]
RL 0.1 $0.913({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.897({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.975({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.041})$ $0.980({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.033})$ $0.936({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.916({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.572$
[…]
GA 0.1 $0.990({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.109})$ $0.905({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.990({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.056})$ $0.928({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $0.998({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.973({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.028})$ $0.062$
[…]
Teacher 0.1 $0.967({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$ $0.898({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.963({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.929({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.018})$ $0.998({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.969({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $0.591$
[…]
FF 0.1 $0.933({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.899({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $7.129({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.195})$ $6.566({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.619})$ $0.131({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.812})$ $0.137({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.808})$ $0.998$
[…]
SSD 0.1 $0.987({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.106})$ $0.902({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.990({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.056})$ $0.926({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $0.998({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.973({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.028})$ $0.063$
[…]
NegGrad+ 0.1 $0.895({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.014})$ $0.898({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.964({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.030})$ $0.950({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.928({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.946({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.044$
[…]
Salun 0.1 $0.936({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.896({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.956({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.022})$ $0.954({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.979({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.939({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.489$
[…]
@@
As for calibration data, it should be independent of train and test data. One requirement for the choice of the calibration set size is that the distribution of the calibration set should match the distribution of the test data as closely as possible. That means a small sample size may lead to distributional dissimilarity and unstable coverage estimation. Therefore, a sufficient sample size of calibration data must be ensured to obtain stable estimates. Figure 1 illustrates the stability of $\hat{q}$ across varying calibration set sizes. We implement them on CIFAR-10 with ResNet-18 in $10\%$ and $50\%$ random data forgetting scenarios. The results shown in Figure 1 are smoothed using B-spline. It reveals that $2,000$ is a sufficient calibration set size for different settings on CIFAR-10 with ResNet-18 to get a stable threshold $\hat{q}$. We also analyze the calibration set size choice of the class-wise forgetting scenario and find that $50$ data points per class as a calibration set is enough. The reason is that, the model’s prediction distribution becomes more complex in a dataset with more classes, requiring a larger calibration set to stabilize the estimation.
2.4 Evaluating Unlearning Methods with New Metrics
Having established the new metrics CR and MIACR, a natural question to ask is:
(Q3) Do existing machine unlearning methods truly achieve forgetting, and how do they perform under our new evaluation metrics?
In this section, we assess the performance of various unlearning methods using the newly introduced metrics, specifically CR, together with coverage and set size. We employ 9 different unlearning methods, RT, FT (Warnecke et al., 2021), RL (Graves et al., 2021), Gradient Ascent (GA) (Thudi et al., 2022), Teacher (Tarun et al., 2023), FisherForgetting (FF) (Golatkar et al., 2020a), SSD (Foster et al., 2024), NegGrad+ (Kurmanji et al., 2024) and Salun (Fan et al., 2023).
The experimental results are presented in Table 4, which summarizes the unlearning performance under $10\%$ random data forgetting scenario on CIFAR-10 with ResNet-18. For more experimental results on CR and MIACR, please check Section 4 and Appendix B. The table 4 includes coverage percentages, set size, CR for both $\mathcal{D}_{f}$ and $\mathcal{D}_{test}$, as well as $\hat{q}$ in different settings of $\alpha$. As observed, it shows a decreasing coverage and set size as the alpha increases in all methods since the threshold $\hat{q}$ is relaxed due to the reduced confidence level.
Based on evaluation criterion ➊, while NegGrad+ outperforms RL under the traditional UA metric, RL exhibits superior performance under the CR metric. This observation suggests that methods excelling in traditional UA metric may not perform well under the CR metric. The underlying rationale behind this is that the CR metric takes into account the possibility that the true labels of some misclassified forget data points may still remain within the prediction set. This observation aligns with the insights we discussed in Section 2.2 regarding the limitation of UA. Based on evaluation criterion ➋, NegGrad+ and RL exhibit the best trade-off between forgetting quality and model performance. FF achieves the lowest CR, indicating the highest forgetting quality, but it exhibits the worst prediction performance on $\mathcal{D}_{test}$. On the other hand, GA, Teacher, SSD, FT and Salun maintain relatively high CR of the forget data $\mathcal{D}_{f}$ across different alpha settings, reflecting their poor forgetting quality.
@@ 3 Enhancing Machine Unlearning via Conformal Prediction
Our analysis indicates that machine unlearning methods generally do not exhibit strong performance under our newly developed evaluation metrics. Even methods with relatively low CR values demonstrate significant compromises in other metrics. This observation raises an important question:
@@
(Q4) Can we explore advanced learning techniques via conformal prediction to optimize unlearning model’s performance under our evaluation metrics?
A training-based machine unlearning method is typically optimized for prediction loss but cannot efficiently support the improvement of forgetting quality in our new metrics. Therefore, we propose a novel and general unlearning framework for training-based machine unlearning methods to enhance machine unlearning. The key innovation of this framework lies in leveraging conformal prediction principles combined with a loss function inspired by the C&W attack (Carlini & Wagner, 2017) to improve forgetting quality.
@@
The original C&W loss function is designed to find adversarial examples that cause misclassification. Here we extend it to an unlearning loss. For forget data $\mathcal{D}_{f}$, the objective of the unlearning loss function is to decrease the model’s confidence in the true labels of $\mathcal{D}_{f}$. Based on this, the C&W-inspired unlearning loss can be defined as:
${\mathcal{L}}_{\text{cw}}({\bm{x}},y_{t})=\max\left(p_{t}({\bm{x}})-\max_{i\neq t}\{p_{i}({\bm{x}})\},\ -\Delta\right),$ (7)
where $({\bm{x}},y_{t})\in\mathcal{D}_{f}$. $p_{i}({\bm{x}})$ is the possibility (after softmax function) of class $y_{i}$, while $z_{t}({\bm{x}})$ is the true label $y_{t}$’s possibility output. We denote $\max_{i\neq t}\{p_{i}({\bm{x}})\}$ as the highest possibility value of the incorrect class. $\Delta$ is the margin parameter (also known as the confidence parameter) that controls the attack strength.
This loss is minimized when $p_{t}({\bm{x}})-\max_{i\neq t}\{p_{i}({\bm{x}})\}\leq-\Delta$ to make $\max_{i\neq t}\{p_{i}({\bm{x}})\}-p_{t}({\bm{x}})\geq\Delta$. It means that the loss maximizes the difference between the highest possibility value for class $y_{i}$ ($i\neq t$) and the possibility value for the true label $y_{t}$, which makes it more difficult for the model to include the true label in the prediction set during conformal prediction process. Additionally, increasing the value of $\Delta$ further reduces the confidence of the true label $y_{t}$ being included in the prediction set. In our work, we set $\Delta=0.01$.
@@
Based on the insight discussed in Section 2, we further improve the C&W-inspired unlearning loss function by combining conformal prediction. In conformal prediction, calibration data helps in estimating non-conformity scores and determining a threshold to ensure valid statistical guarantees about the model’s uncertainty estimates. Therefore, we reserve a portion of calibration data $\mathcal{D}_{c}^{\prime}$ for the unlearning phase, which is kept separate from the calibration data $\mathcal{D}_{c}$ used in the evaluation phase.
According to Eq. 1, calibration data $\mathcal{D}_{c}^{\prime}$ is used to calculate non-conformity scores and generate a threshold $\bar{q}$ based on an $\alpha$. Then, we calculate the non-conformity scores of $\mathcal{D}_{f}$ to obtain the corresponding prediction set. Then, by revising C&W-inspired unlearning loss with this calibration step, a general unlearning loss function is defined as follows:
${\mathcal{L}}_{\text{unlearn}}({\bm{x}},y_{t})=\max\left(S({\bm{x}},y_{t})-\bar{q},\ -\Delta\right),$ (8)
where $({\bm{x}},y_{t})\in\mathcal{D}_{f}$. We replace $p_{t}({\bm{x}})$ in Eq. 7 with the non-conformity score $S({\bm{x}},y_{t})$ of true label $y_{t}$, and replace $\max_{i\neq t}\{p_{i}({\bm{x}})\}$ with $\bar{q}$ which will be updated before each epoch. It ensures that ${\mathcal{L}}_{\text{unlearn}}$ adheres to the principle of ${\mathcal{L}}_{\text{cw}}$, which encourages $\bar{q}-S({\bm{x}},y_{t})\geq\Delta$.
@@
It is a general framework applicable across various training-based machine unlearning methods. To preserve the effectiveness of specific machine unlearning methods, we retain their original loss ${\mathcal{L}}_{\text{original}}$. Consequently, we sum these terms to form the final objective loss function as follows:

### v2

@@ 2.3.2 Discussion of Confidence Level and Calibration Set Size
In conformal prediction, the confidence level $1-\alpha$ (i.e., miscoverage rate $\alpha$) and calibration set size are two crucial factors for prediction set coverage and reliability. We next discuss the suitable settings for the confidence level and calibration set size, and the rationale behind them.
Figure 2: The stability of $\hat{q}$ in different calibration set sizes. When the calibration set size is greater than $2000$, the fluctuations of $\hat{q}$ remain within a stable range.
Confidence Level $\bm{1-\alpha}$.
A smaller miscoverage rate $\alpha$, i.e., a higher confidence level $1-\alpha$, guarantees more reliable coverage, while a higher value of $\alpha$ results in a lower confidence level. In the conformal prediction related works Angelopoulos and Bates (2021); Papadopoulos et al. (2002); Romano et al. (2020); Tailor et al. , $\alpha=0.05$ has become the standard in most cases. This value can be rooted in its widespread adoption in statistical hypothesis testing, where it balances the trade-off between false positives and practical usability. Following previous work, we set $\alpha$ to $0.05$, but we still report results for a higher range of $\alpha$ values ($0.10$, $0.15$, and $0.20$) in Appendix B.3 and B.4 to account for scenarios where a more relaxed confidence level may be needed. Unless otherwise stated, all subsequent analyzes use the recommended default of $\alpha=0.05$.
Calibration Set Size.
@@
As for calibration data, a portion of the validation data is set aside as calibration data, ensuring it remains independent from both the training and test data. A key requirement in selecting the calibration set size is to avoid producing an abnormal threshold $\hat{q}$, as a small sample size may introduce outliers and lead to unstable coverage estimates. Therefore, it is essential to ensure a sufficiently calibration set to obtain reliable and stable estimates. Figure 2 illustrates the stability of $\hat{q}$ across varying calibration set sizes. The results are smoothed using B-spline. We implement them on CIFAR-10 with ResNet-18 in $10\%$ and $50\%$ random data forgetting scenarios. The results show that for different settings using ResNet-18 on CIFAR-10, after the calibration set size is larger than $1000$, abnormal $\hat{q}$ values do not occur anymore, and a stable threshold $\hat{q}$ can be obtained. Similarly, we analyze the calibration set size of the class-wise forgetting scenario and find that fewer calibration data points are required compared to random data forgetting. This is because the targeted removal of specific classes reduces the complexity of the distribution, unlike the broader variability introduced by random data removal.
@@ 3 Enhancing Machine Unlearning via Conformal Prediction
Based on the findings in Table 2 in Section 2.1 and Table 6 in Appendix B.1, we observe that existing training-based unlearning methods are typically optimized with respect to loss functions that do not directly support the improvement of forgetting quality from our fake unlearning perspective. Specifically, the optimization objectives of existing methods fail to ensure that the ground truth labels are sufficiently pushed out of the conformal prediction set, which is key to overcoming fake unlearning. This raises a critical question:
@@
(Q4) Can we explore advanced unlearning techniques via conformal prediction to optimize existing unlearning model’s forgetting quality?
Therefore, we propose a novel and general conformal prediction-based unlearning framework (CPU) tailored for training-based unlearning methods, aimed at enhancing their forgetting quality. A key insight driving our framework is to overcome the issue exposed by fake unlearning. This emphasizes that the non-conformity scores of ground truth labels should be pushed as far beyond the conformal prediction threshold $\hat{q}$ as possible. Interestingly, this goal aligns naturally with the design of the C&W attack loss Carlini and Wagner (2017), which motivates our creative adaptation to the unlearning scenario.
@@
Let us first apply the original C&W loss directly to the unlearning scenario, without yet incorporating conformal prediction. For forget data $\mathcal{D}_{f}$, the goal of the unlearning loss is to decrease the model’s confidence in the true labels of $\mathcal{D}_{f}$. Based on this, the C&W-inspired unlearning loss is defined as:
${\mathcal{L}}_{\text{cw}}({\bm{x}},y_{t})=\max\{p_{t}({\bm{x}})-\max_{i\neq t}\{p_{i}({\bm{x}})\},\ -\Delta\},$ (8)
where $({\bm{x}},y_{t})\in\mathcal{D}_{f}$ and ${}\max\{\cdot\}$ is a maximum operator that selects the largest value from the set. $p_{i}({\bm{x}})$ is the probability of class $y_{i}$, and $p_{t}({\bm{x}})$ refers specifically to the probability assigned to the true label $y_{t}$. We denote $\max_{i\neq t}\{p_{i}({\bm{x}})\}$ as the highest probability value of the non-ground truth classes. This loss ${\mathcal{L}}_{\text{cw}}$ maximizes the difference between the highest probability value for class $y_{i}\ (i\neq t)$ and the probability value for the true class $y_{t}$. It tries to decrease the probability of true class $y_{t}$ and further increase that of the class $y_{i}$ with the highest probability. The margin parameter $\Delta$ controls the enforced margin between the true class and the strongest competing class. When the $\max_{i\neq t}\{p_{i}({\bm{x}})\}-p_{t}({\bm{x}})<\Delta$, this loss encourages the model to decrease the true label’s probability $p_{t}({\bm{x}})$. Increasing the value of $\Delta$ further increase the margin between $\max_{i\neq t}\{p_{i}({\bm{x}})\}$ and $p_{t}({\bm{x}})$.
With this C&W loss, we can indeed reduce the probability assigned to the true label $y_{t}$, thereby compelling the model to misclassify the data point into another class $y_{i}$. However, this loss still fails to guarantee that the true label $y_{t}$ can be excluded from the conformal prediction set. If we let the threshold in conformal prediction play the role of $\max_{i\neq t}\{p_{i}({\bm{x}})\}$ in Eq. 8, and push non-conformity score of $y_{t}$ further away from this threshold, the above issue can be effectively resolved. Therefore, we further improve the C&W-inspired unlearning loss function by combining conformal prediction.
@@
In conformal prediction, calibration data helps in estimating non-conformity scores and determining a threshold to ensure valid statistical guarantees about the model’s uncertainty estimates. A portion of calibration data $\mathcal{D}_{c}^{\prime}$ can be reserved for the unlearning phase, which is kept separate from the calibration data $\mathcal{D}_{c}$ used in the evaluation phase. With calibration data $\mathcal{D}_{c}^{\prime}$, the threshold $\bar{q}$ for the unlearning phase is easily calculated given an $\alpha$. Given $\bar{q}$, by revising C&W-inspired unlearning loss with a calibration step, a general unlearning loss function is defined as follows:
${\mathcal{L}}_{\text{unlearn}}({\bm{x}},y_{t})=\max\{\bar{q}-S({\bm{x}},y_{t}),\ -\Delta\}.$ (9)
We replace probability $p_{t}({\bm{x}})$ and $\max_{i\neq t}\{p_{i}({\bm{x}})\}$ in Eq. 8 with the threshold $\bar{q}$ and non-conformity score $S({\bm{x}},y_{t})$ respectively. The threshold $\bar{q}$ is updated in each training epoch to obtain an accurate value. Since $\bar{q}$ is computed merely as a quantile, this process incurs negligible computational overhead.
The loss ${\mathcal{L}}_{\text{unlearn}}$ adheres to the same principle of ${\mathcal{L}}_{\text{cw}}$, which encourages $S({\bm{x}},y_{t})-\bar{q}\geq\Delta$. It helps to increase non-conformity score $S({\bm{x}},y_{t})$ of true label $y_{t}$ to surpass the threshold $\bar{q}$. As an improvement over the loss ${\mathcal{L}}_{\text{cw}}$, the loss ${\mathcal{L}}_{\text{unlearn}}$ makes it more difficult for the model to include the true label in conformal prediction set. In this loss, even a small value of $\Delta$ is sufficient to achieve the desired effect, because the true label $y_{t}$ is excluded from the conformal prediction set once its non-conformity score $S({\bm{x}},y_{t})$ exceeds the threshold $\bar{q}$. Therefore, in our work, we set $\Delta=0.01$.
@@
As a general framework, to preserve the efficacy of specific unlearning methods themselves, we reserve their original loss ${\mathcal{L}}_{\text{original}}$ in our framework. Consequently, we combine these terms to form the final objective loss function as: